### v2

# {cutout}02.753cm0cm1 empty ViC-Bench     : Benchmarking Visual-Interleaved Chain-of-Thought Capability in MLLMs with Free-Style Intermediate State Representations

Abstract
Visual-Interleaved Chain-of-Thought (VI-CoT) enables MLLMs to continually update their understanding and decisions based on step-wise intermediate visual states (IVS), much like a human would, which demonstrates impressive success in various tasks, thereby leading to emerged advancements in related benchmarks. Despite promising progress, current benchmarks provide models with relatively fixed IVS, rather than free-style IVS, whch might forcibly distort the original thinking trajectories, failing to evaluate their intrinsic reasoning capabilities. More importantly, existing benchmarks neglect to systematically explore the impact factors that IVS would impart to untamed reasoning performance. To tackle above gaps, we introduce a specialized benchmark termed ViC-Bench, consisting of four representive tasks: maze navigation, jigsaw puzzle, embodied long-horizon planning, and complex counting, where each task has dedicated free-style IVS generation pipeline supporting function calls. To systematically examine VI-CoT capability, we propose a thorough evaluation suite incorporating a progressive three-stage strategy with targeted new metrics. Besides, we establish Incremental Prompting Information Injection (IPII) strategy to ablatively explore the prompting factors for VI-CoT. We extensively conduct evaluations for 18 advanced MLLMs, revealing key insights into their VI-CoT capability. Our proposed benchmark is publicly open at [Uncaptioned image] Huggingface.
11footnotetext: Work done during research internship at Meituan-M17.22footnotetext: Equal Contribution.33footnotetext: Project Lead.
1 Introduction
AI field is currently evolving from LLMs [45, 9, 69] to MLLMs [8, 26, 74, 72], which integrate various modalities into the language decoders [73]. Achieving human-level multi-modal intelligence requires transcending basic perceptual capabilities to attain sophisticated reasoning. Drawing inspirations from the remarkable success of Chain-of-Thought (CoT) in LLMs [11, 38, 24], the integration of CoT into visual-language contexts has catalyzed transformative progress, giving rise to visual CoT [33, 73].
The initial visual CoT involves vision signals only as input, whereas the entire rationales are composed of language, in which various methods and related benchmarks make rapid advancements [10, 6, 77]. However, this paradigm overlooks the explicit visual representation updates and continuous understanding of visual feedbacks, misaligning with the human cognitive process of using visual thoughts for concrete reasoning and textual thoughts for abstract reasoning. To this end, Visual-Interleaved Chain-of-Thought (VI-CoT), which incorporates step-wise intermediate visual states (IVS) based on visual inputs, has made rapid progress. According to the source of IVS, current VI-CoT methods are primarily divided into two types. The first type involves autonomously generating IVS based on its internalized understanding [76, 33, 84]. However, this approach currently struggles due to the limited generative capabilities of MLLMs [76, 84]. The second type involves providing IVS through external knowledge retrieval, utilizing expert tools or human in the agent-form, which shows impressive results in broad tasks [81, 13, 34]. Meanwhile, to evaluate the developments of recent VI-CoT methods, various benchmarks have emerged [13, 82]. Despite promising advancements, few of them provides free-style IVS representations to MLLMs, as illustrated in Tab. 1. CoMT [13] primarily provides fixed IVS, which might forcibly distort the original planning trajectories. While MageBench [82] offers the dynamic IVS but imposes the attribute constraints of action-observation memory. More importantly, existing benchmarks [60, 28, 13, 6] neglect to systematically assess the impact factors that IVS would impart to untamed reasoning performance in MLLMs. (i.e., Positive, Negative, or Null). As a result, a natural question arise: Could MLLMs leverage VI-CoT, which closely aligns with human cognitive behavior, to inherently achieve better reasoning performance?
Figure 1: Performance overview of advanced MLLMs on four tasks in average ACC across stages.
To tackle above gaps, we introduce the specialized ViC-Bench for evaluating VI-CoT, selecting four representative tasks (i.e., maze navigation, jigsaw puzzle, embodied long-horizon planning, and complex counting), which require models to dynamically interact with visual contexts and continuously update their understanding and decision-making based on step-wise IVS. We first propose dedicated data construction pipeline, resulting in 250 unique images for each task. We then introduce free-style IVS generation workflows supporting function calls to sufficiently support the investigation on VI-CoT performance. Based on the constructed data, we propose a novel evaluation suite incorporating progressive three-stage evaluation strategy with targeted new metrics. Specifically, Stage 1 involves multiple-choice QA, while Stage 2 focuses on open-ended QA, both utilizing visual signals solely at the input to establish solid foundations for Stage 3. Building on this, Stage 3 features open-ended QA with free-form IVS, enriching from the hierarchical references for VI-CoT evaluation through our progressive design. Subsequently, we define new Recall and ThinkGain metrics by black-boxing the retrospection process, along with a Legality metric to tackle the clear rule boundaries, thereby establishing a comprehensive suite for each stage. To ablatively explore the prompting factors affecting VI-CoT capability, we further design the Incremental Prompting Information Injection (IPII) strategy across three stages, utilizing varying global-aware prompting levels. We totally examine 18 SOTA MLLMs on our ViC-Bench, offering extensive quantitative and qualitative results, as illustrated in Fig. 1. We uncover significant performance gaps between open-source and proprietary MLLMs in both quantitative analyses and prompting studies. Moreover, most MLLMs show substantial disparities compared to human-level proficiency. For example, powerful Gemini-2.5-pro [63] performs inadequately on jigsaw puzzle, only achieving 6.80% ACC in Stage 2.
In addition to providing insights into VI-CoT capabilities of MLLMs, ViC-Bench can also establish foundations for the developments of native MLLMs [59, 61], multi-modal agents [79, 56], embodied AI [34, 35], and autonomous driving [15]. The main contributions of this paper are three-fold:
  • $\bullet$
    We curate ViC-Bench, including four representative tasks, each with dedicated construction and free-style IVS generation pipelines. Moreover, we engage human-machine collaborations in both construction and evaluation to deliver a high-quality benchmark for VI-CoT reasoning.
  • $\bullet$
    We propose a thorough evaluation suite that includes a progressive three-stage evaluation strategy with newly targeted metrics to meticulously examine inherent VI-CoT performance using free-style IVS. We further introduce the IPII strategy to ablatively explore the prompting factors for VI-CoT.
  • $\bullet$
    Extensive experiments and analyses are conducted on 18 proprietary and open-source MLLMs. We summarize several key observations and insights, hoping to inspire advancements of future works.
Table 1: The statistics comparisons of ViC-Bench and related representative benchmarks. #No.: Unique sample number. Free: Free-style IVS supporting function calls.
Benchmark Venue Source #No. Task Multi-Step IVS Free-Style
M3CoT [10] (Test Part) ACL’24 Web 2,358 3 $\times$ $\times$
LEGO-Puzzles [60] arXiv’25 Synthesized 1,100 3 $\times$ $\times$
MME-CoT [28] ICML’25 Web 808 6 $\times$ $\times$
CoMT [13] AAAI’25 Web 3,853 4 $\times$
MageBench [82] arXiv’24 Synthesized & Web 483 3
VERIFY [6] arXiv’25 Web 600 1 $\times$ $\times$ $\times$
ViC-Bench (Ours) Synthesized & Web 2,751 4
2 Related Work
Multi-modal Large Language Models. The integration of multi-modal information with LLMs [45, 9] leads to the emergence of MLLMs [8, 26, 74, 66], exhibiting impressive performance in various multi-modal understanding and generation tasks. MLLMs can be broadly categorized into two types: pipeline-based and native paradigms. The pipeline-based MLLMs can be classified into three types based on the multi-modal integration strategies: (1) Feature mapping with MLPs, such as PaLM-E [19], LLaVA [39], and CogVLM [71]; (2) Query-based cross-attention components (e.g., InstructBLIP [16], Mini-GPT4 [86], Qwen-VL series [4, 70, 5]); (3) Cross-attention layers within LLMs, such as Flamingo [1] and IDEFICS [30]. Meanwhile, MLLMs integrated with generation capabilities through coupled components have also been largely promoted [51, 59, 85]. As for native MLLMs, they primarily achieve unified understanding and generation through auto-regressive manners with elaborate tokenizers [75, 40, 61]. Most recently, the release of OpenAI o3/o4 [49] and DeepSeek-R1 [24] sparks a wave of interests in reasoning enhancements, highlighting the effectiveness of CoT [35, 80]. Inspired by this, researchers advance reasoning capabilities of MLLMs employing visual CoT mechanisms [43, 55, 41].
Visual-Interleaved Chain-of-Thought. VI-CoT involves the engagement of step-wise IVS through the entire reasoning process, achieving impressive performance across various downstream scenarios. However, due to limited generative capabilities, MLLMs struggle to generate the native IVS, which are essential for the in-context knowledge retrieval. As a result, current methods primarily rely on external knowledge retrieval to develop IVS, such as expert tools or human in the agent-form [73, 38]. CMMCoT [81] utilizes the visual region tokens as supervisory signals to perform interleaved reasoning. Zhang et al. [83] extend o1-style reasoning to interactive embodied search. Gao et al. [23] propose the attention-driven selection method to realize interleaved CoT. VoT [21] breaks down complex task into sub-problems, and address them from low to high using scene graphs. MVoT [33] enables visual thinking by generating visual visualizations of reasoning traces. Hu et al. [25] provide MLLMs with sketchpad and expert tools to conduct interleaved CoT. Meanwhile, related benchmarks have emerged to extensively evaluate various VI-CoT methods. CoMT [13] constructs four types of visual operations, requiring multi-modal reasoning outputs. Zhang et al. [82] propose MageBench for evaluating the MLLMs’s capabilities of being an agent. [83] cultivates 809 test cases across 12 scenarios for hierarchical long-horizon tasks. Despite great advancements, few of these benchmarks provide the free-form IVS representations and systematically evaluate the impacts that IVS might have on the untamed reasoning performance. To bridge these gaps, we carefully construct four representative VI-CoT tasks with the free-form IVS representations and further propose a comprehensive evaluation suite incorporating three stages.
3 ViC-Bench Construction
3.1 Overview
We introduce ViC-Bench, consisting of four representative VI-CoT tasks: Maze Navigation (Sec. 3.2), Jigsaw Puzzle (Sec. 3.3), Embodied Long-Horizon Planning (Sec. 3.4), and Complex Counting (Sec. 3.5). The overall construction workflow can be mainly divided into Raw Data & Pre-processing, Three-Stage Construction, IVS Generation, and Human Recheck, as displayed in Fig. 2.
3.2 Maze Navigation
Raw Data & Pre-processing. As shown in Fig. 2 (a), we utilize Maze [27] library coupled with DFS method to render 4 $\times$ 4 mazes in multiple batches. After generation, we first screen out those with navigation lengths ranging from 5-8, then aggregate mazes with the same starting point. Finally, mazes with duplicate shortest path after global-aware aggregation are removed to ensure uniqueness.
Stage 1. For the processed mazes, we randomly select an original maze $\mathbf{M_{1}}$ in the non-overlapping paradigm, then select three other original mazes $\mathbf{M_{2}}$, $\mathbf{M_{3}}$, and $\mathbf{M_{4}}$. Subsequently, we draw the endpoints of $\mathbf{M_{2}}$-$\mathbf{M_{4}}$ onto $\mathbf{M_{1}}$ to establish three incorrect options, facilitating MLLMs to select the correct endpoint based on the starting point and the given navigation path.
Stages 2 & 3. Based on $\mathbf{M_{1}}$ of Stage 1, we mark the starting and corresponding endpoints with $\mathbf{S}$ and $\mathbf{E}$, and require models to respond with the correct path from $\mathbf{S}$ to $\mathbf{E}$ under the clear rules. Stage 3 further builds upon Stage 2, with its main feature being the application of free-style IVS in response to the step-wise instructions of models under the agent-form, promoting to investigate the untamed VI-CoT boundaries.
IVS Generation. Based on simulated functions, we perform multi-step simulations on the input maze, employing a blue pentagram to mark the agent position. We set the maximum attempts to 30.
Human Recheck. Throughout three-stage data construction, we employ human experts to perform one-by-one recheck on the 250 mazes to sufficiently ensure data feasibility. Besides, we conduct manual quality inspections on the meta outcomes of Stage 3 to faciliate the stability of our evaluations.
3.3 Jigsaw Puzzle
Raw Data & Pre-processing. To eliminate the risk of data leakage from the familiar datasets [17, 36], we construct an image source pool using elaborate prompts with DALLE-3 [46], FLUX.1-schnell [29], Kwai-Kolors [65], Stable-Diffusion-3 Medium [52], WanX [2], and Midjourney-V6.1 [44], which can sufficiently ensure the image diversity and uniqueness. Following [31, 37], we employ the manual white-box approach to filter the generated images, primarily considering three discriminative metrics: T2I consistency, reasonableness, as well as realism, i.e.,
$S_{overall}=0.6\times S_{consistency}+0.2\times S_{reasonableness}+0.2\times S_{realism},$ (1)
where $S\in[1,5]\cap\mathbb{Z}$. Based on the descending order of overall scores, we finally select 250 images.
Stage 1. We first adjust the selected images to 224 $\times$ 224 resolutions and divide them into 4 $\times$ 4 patches. We then targetedly extract six patches using proposed weighted dispersed sampling strategy. Afterwards, the selected patches are removed from the original image to generate masked image $\mathbf{I}_{m}$. Subsequently, the chosen patches are randomly numbered and concatenated with $\mathbf{I}_{m}$ along the vertical direction to output the overall input image, as illustrated in Fig. 2 (b). As for options, four patches are first randomly selected and correctly positioned to establish the correct option. Three incorrect options are then generated based on the puzzle states with either (3 ✓& 1 $\times$) or (2 ✓& 2 $\times$) patch placements, leading to the plausible yet incorrect paradigm, which challenge the discriminative capabilities of MLLMs in region-aware semantic understanding and spatial-aware reasoning.
Stages 2 & 3. The overall input images of last two evaluation stages almost keep the same with Stage 1, but additionally label the six vacant regions in $\mathbf{I}_{m}$. Moreover, the arrangement of six patches in the overall input image underneath remains consistent with Stage 1. The task instructions here require models to directly output the correct mapping between patches and vacant regions.
IVS Generation. Based on the simulated functions and planning trajectories of MLLMs, we conduct multi-step agent-form evaluations with free-style IVS. Besides, we establish the rule boundaries such that if a patch or vacant region is repeatedly utilized, causing conflicts, we will deem the current action invalid and provide appropriate guidance. We also set the maximum attempts to 30.
Human Recheck. Following maze navigation, we also employ human experts to recheck the generated puzzles across stages and perform quality inspections on VI-CoT evaluations in Stage 3.
Figure 2: The illustrations of the overall construction pipelines for four downstream tasks.
3.4 Embodied Long-Horizon Planning
Raw Data & Pre-processing. We establish this task based on the benchmark from Embodied-reasoner [83]. We first consider the length of action sequences for manual selection (i.e., longer than 20 steps are removed) and further ensure the uniqueness of each sequence. Semantic aggregations are then performed to promote sequence separation. In the end, we manually select 250 samples.
Stage 1. We regard the final visual state achieved after completing the sequence as the correct option. Three incorrect options are generated using the ambiguous IVS, which are extracted from the original trace and manually refined to increase difficulty, as displayed in Fig. 2 (c). This setup thoroughly tests the long-horizon planning capabilities of models, providing baselines for latter evaluations.
Stage 2. For each sample, we take the initial observation of action sequence in Stage 1 as the starting state $\mathbf{O}_{s}$, and the completely achieved state as the ending state $\mathbf{O}_{e}$. MLLMs in this stage are required to open-endedly respond with the action sequence which can successfully transition from $\mathbf{O}_{s}$ to $\mathbf{O}_{e}$.
Stage 3 & IVS Generation. Based on the construction priors in Stages 1 & 2, we employ human experts to perform Stage 3 assessment in the real-world scenarios. In this way, we deploy the human-machine-in-the-loop paradigm, where MLLM as a planner and human as an actioner [82], to conduct free-style IVS enhanced evaluations. Concretely, human experts capture the intermediate images based on the model’s thinking plans and upload them to the cloud storage via a dedicated mobile APP, subsequently utilizing function calls for the step-wise collaborative reasoning evaluation.
3.5 Complex Counting
Raw Data & Pre-processing. We build this task based on the samples from JHU-CROWD++ [57] and CoMT [13]. According to annotations in [13], we first filter out samples with headcounts ranging from 10 to 40, then remove those that are ambiguous or difficult to delineate. As displayed in Fig. 2 (d), we then manually draw irregular lines to separate the input image into four regions labeled as $\mathbf{\{1,2,3,4\}}$, leading to the completed input image $\mathbf{H}$. In the end, we totally construct 250 samples.
Stage 1. We manually count the visible heads and regard the correct headcounts as the ground truth option, tackling limitations in previous work [13] that solely relies on DNNs for annotations. We then employ human experts to establish three incorrect options, leading to higher selection difficulties.
Stages 2 & 3. The input format of image remains consistent with Stage 1, but the QA format changes from multiple-choice to open-ended, where we require models to directly respond with the headcounts for each region. The output format is defined as <Begin> ($\mathbf{1}$, $\mathbf{h_{1}}$), ($\mathbf{2}$, $\mathbf{h_{2}}$), ($\mathbf{3}$, $\mathbf{h_{3}}$), ($\mathbf{4}$, $\mathbf{h_{4}}$) </End>, where $\mathbf{h_{i}}$ ($\mathbf{i}\in\{1,2,3,4\}$) denotes the headcount in each separated region.
IVS Generation. Considering that complex counting does not hold the explicit dynamics of action sequences like maze navigation or jigsaw puzzle, and the empirical experimental results further indicate that overly fine-grained masks can lead to overthinking and object hallucinations, we thus utilize the coarse-grained region-aware mask instead of fine-grained single head mask for the IVS generation. As a result, we manually draw bounding boxes for $\mathbf{H}$ to accurately cover masks over all the heads in each separated region, consistently leading to four IVS for each data sample.
Figure 3: The overall illustration of our evaluation workflow taking maze navigation as the example.
4 Evaluation Suite
4.1 Progressive Three-Stage Evaluation Strategy
As illustrated in Fig. 3, we propose a holistic progressive three-stage evaluation strategy, rather than simply focusing on the final answer, to examine the untamed understanding of VI-CoT capability in MLLMs. Stage 1 only focus on the final results, employing the common multiple-choices QA [53, 18, 42] to preliminarily determine the visual CoT performance. Based on POMDP [82], Stage 1 can be formulated as:
$\pi_{\theta}(p_{sys},(Q,\mathbf{v_{0}},C),p_{task},p_{cot},p_{io})\rightarrow(\mathbf{r_{1}},\mathbf{r_{2}},\mathbf{r_{3}},...,\mathbf{r_{T}},\mathbf{C_{S}}),$ (2)
where $p_{sys}$ and $p_{task}$ denote system and task prompts. $p_{cot}$ and $p_{io}$ refer to CoT and IO prompts, designating the inner thought flow and output format. $(Q,\mathbf{v_{0}},C)$ represents question and inital observation with options. $\mathbf{r_{i}}$ ($\mathbf{i}\in\{1,...,\mathbf{T}\}$) are the reasoning rationales. $\mathbf{C_{S}}$ is the selection. In Stage 2, we convert the multiple-choices QA into the more challenging open-ended format, i.e.,
$\pi_{\theta}(p_{sys},(Q,\mathbf{v_{0}}),p_{task},p_{cot},p_{io})\rightarrow(\mathbf{r_{1}},\mathbf{r_{2}},\mathbf{r_{3}},...,\mathbf{r_{T}},\mathbf{A_{F}}),$ (3)
where $\mathbf{A_{F}}$ is the formatted answer. Doing so allows us to observe the reasoning performance in an open-ended manner, which leads to direct subjective Answer-Only evaluations, thereby serving as an important indicator for Stage 3. Finally, Stage 3 pays attention to the legal free-style evaluations using function calls in the agent-form to explore in-depth thinking gains brought by the IVS representations, addressing the shortcomings in previous works where fixed IVS might forcefully influence the inherent planning in MLLMs by constraining the judgment path. The overall process can be represented as:
$F(p_{sys},(Q,\mathbf{v_{0}}),p_{task},p_{cot},p_{io};\pi_{\theta})\rightarrow\mathbf{R},$ (4)
$\pi_{\theta}(p_{sys},(Q,\mathbf{v_{0}}),((\mathbf{a_{1}},\mathbf{v_{1}}),...,(\mathbf{a_{t}},\mathbf{v_{t}})),p_{task},p_{cot},p_{io})\rightarrow(\mathbf{a_{t+1}},\mathbf{v_{t+1}}),$ (5)
where $\mathbf{R}$ denotes the final answer and $(\mathbf{a_{t}},\mathbf{v_{t}})$ refers to action and IVS feedback at the $t^{th}$ step. Our free-style exploration in the agent-form further stimulates the impact of IVS, thereby more comprehensively investigating the untamed VI-CoT capability inherented in current MLLMs.
4.2 Evaluation Metrics
As shown in Fig. 3, Stage 1 employs Accuracy (ACC) metric, Stage 2 includes ACC, Recall_o, and Legality metrics, Stage 3 further enhances Stage 2 with the newly introduced ThinkGain metric.
Stage 1. We mainly focus on whether the final choice is correct or not, consistently utilizing ACC as the sole metric across tasks. As shown in Fig. 3 (a), we first perform rule-based extraction to obtain the selected options, then directly compare them with Ground Truth to compute the ACC metric, i.e.,
$ACC=\frac{TP+TN}{TP+TN+FP+FN},$ (6)
where $TP$, $TN$, $FP$, and $FN$ denote the number of true positives, true negatives, false positives, and false negatives. We can tentatively observe reasoning performance, offering extensive references.
Stage 2. ACC only focuses on the correctness of final answer but neglects how many informative steps the model has taken to reach the correct answer. However, examining the rigorousness of models towards Ground Truth is very crucial, as it can delineate the response progress and provide insights into the effectiveness of open-ended visual CoT in MLLMs. To tackle above gaps, we introduce task-specific Recall_overall metric (i.e., $\mathbf{R\_o}$), only considering the final output. $\mathbf{R\_o}$ measures the informative steps the model has achieved to emphatically characterize its reasoning ability, i.e.,
[…]
where $\mathcal{S}$ denotes the key step-wise components for tasks and $\mathcal{S}^{k}$ refers to the $k^{\text{th}}$ method of a question.
In maze navigation, we extract the final predictions by the rule-based method, then determine the legal endpoint $\mathbf{P_{f}}$ reached by the predicted path using our rendering simulations. We then compare $\mathbf{P_{f}}$ with Ground Truth since each constructed maze has a unique shortest path. If $\mathbf{P_{f}}$ exists in the coordinate set $\mathbf{C_{set}}$ of Ground Truth, we thus determine $\mathcal{S}_{\text{reached}}^{k}$ based on the distance from $\mathbf{P_{f}}$ to the starting point $\mathbf{S}$. If $\mathbf{P_{f}}$ is not in $\mathbf{C_{set}}$, we directly take $\mathbf{R\_o}$ as 0. In jigsaw puzzle, we follow maze navigation to extract final predictions. Considering the uniqueness of puzzle correspondence sequence, we directly perform global-aware strict matching under the component-wise paradigm between predictions and Ground Truth. We calculate $\mathbf{R\_o}$ through taking the number of correctly filled patches as $\mathcal{S}_{\text{reached}}^{k}$ and the counts of extracted patches as $\mathcal{S}^{k}$. Considering the atomic action of embodied long-horizon planning task have dependency constraints, we verify the correctness of atomic action utilizing semantic matching [14] rather than strict matching, keeping consistent with [83]. A pair-wise match is regarded as successful only when the semantic similarity of atomic action pair exceeds 0.95. For ACC, the sample is taken as correct if all the atomic actions are validly matched. For $\mathbf{R\_o}$, we take the accumulated effective length of verified actions and the total predicted steps as $\mathcal{S}_{\text{reached}}^{k}$ and $\mathcal{S}^{k}$, respectively. As for complex counting, we follow jigsaw puzzle to perform global-aware pair-wise matching, except that we extra implement a fault tolerance mechanism due to the high difficulty of this task for current MLLMs, with a tolerance set to 1.
The compliance with explicit rules in maze navigation and jigsaw puzzle poses challenges to the instruction-following and visual perception capabilities of MLLMs, which correlatively promotes VI-CoT evaluation, leading to our Legality metric. In maze navigation, we consider two types of illegal behaviors: going out of bounds and hitting walls. For jigsaw puzzles, the illegal behaviors are: repeated patch placements and repeated filling of vacant regions. Specifically, we segment the predictions followed by conducting step-wise simulation rendering to determine the legal steps, i.e.,
$\text{Legality}=\frac{S_{L}}{S_{O}},$ (8)
where $S_{L}$ and $S_{O}$ refers to the number of legal steps and overall partition steps, respectively.
Stage 3. Building upon Stages 1 & 2, we further introduce a new metric denoted ThinkGain to examine the influence of free-style IVS on VI-CoT reasoning performance. Drawing inspirations from the rewards of GRPO in [24], we black-box the retrospection process and focus only on the decision states ($\mathbf{D_{d}}$ $\mathbf{D_{o}}$) both before the retrospection commences and after it concludes, avoiding the negative impact of numerous ongoing factors on VI-CoT evaluation. We then employ the Recall metric defined in Stage 2 to assess $\mathbf{D_{d}}$ and $\mathbf{D_{o}}$. Overall, the ThinkGain metric can be represented as:
$\text{ThinkGain}=\mathbf{R\_o}-\mathbf{R\_d},$ (9)
where $\mathbf{R\_d}$ denotes the Recall metric calculated with $\mathbf{D_{d}}$. Besides, the definition of $\mathbf{D_{d}}$ varies due to inconsistence in task representations. In maze navigation, we regard $\mathbf{D_{d}}$ as the terminal point reached before the first retrospection. In jigsaw puzzle and complex counting, we treat the state of each patch or region upon its first utilization as $\mathbf{D_{d}}$. For embodied long-horizon planning, we directly employ human experts to determine two decision states.
4.3 Incremental Prompting Information Injection Strategy
Based on three-stage evaluations and metrics, we introduce the Incremental Prompting Information Injection (IPII) strategy to ablatively explore the prompting factors for VI-CoT under three varying levels of global-aware prompting information. Concretely, Level-1 establishes a baseline only using the original prompts of three-stage evaluations. Level-2 involves implicit VI-CoT prompts, which guides models to update its internal step-wise IVS, thereby compelling models to leverage its visual imagination for subsequent planning trajectory, rather than solely relying on the initial observation. Level-3 further augments prompts with external knowledge, providing supporting information to externally enhance the visual perception.
5 Experiments
Table 2: The performace evaluations of advanced MLLMs in our metrics (%) on maze navigation. Note that we highlight the best performance in bold and underline the second performance.
Method Organization Reason- oriented Stage-1 Stage-2 Stage-3
ACC ACC Recall_o Legality ACC Recall_o ThinkGain Legality
Commercial proprietary MLLMs
GPT-4o-20240513 [47] OpenAI $\times$ 50.00 0.00 11.65 53.18 66.40 69.41 33.01 63.76
o3 [48] OpenAI 87.60 17.52 29.11 70.55 74.40 76.86 44.05 81.33
o4-mini [49] OpenAI 94.40 15.45 22.33 68.45 58.40 61.78 27.98 69.88
Gemini-2.5-pro-preview-03-25 [63] Google 94.00 6.80 12.95 55.05 68.80 70.27 32.96 73.28
Gemini-2.0-flash [62] Google 60.80 0.00 11.35 54.94 53.20 59.48 28.11 64.79
Claude-3.5-Sonnet-20241022 [3] Anthropic 53.20 1.60 11.80 52.07 25.60 27.90 6.23 84.63
Step-1o-vision-32k [58] StepFun $\times$ 46.40 0.80 10.72 51.08 16.40 21.64 4.84 30.87
Doubao-1.5-vision-pro-32k [7] ByteDance $\times$ 44.00 1.20 10.89 52.03 13.20 28.53 6.22 58.40
SenseChat-Vision-V5.5 [54] SenseTime $\times$ 62.80 0.00 11.22 53.03 22.40 34.36 10.50 63.92
Hunyuan-Vision-20250103 [68] Tencent $\times$ 30.00 0.40 8.89 48.73 16.40 27.86 7.03 46.57
Open-source MLLMs
LLaVA-OneVision-72B [32] ByteDance & NTU $\times$ 38.00 0.00 8.77 52.25 18.40 30.44 8.17 49.80
InternVL-2.5-78B [12] Shanghai AI Lab $\times$ 46.80 0.00 10.75 53.47 29.60 36.31 11.27 60.52
InternVL-3-78B [87] Shanghai AI Lab $\times$ 43.60 0.00 11.74 52.73 37.60 45.03 12.21 61.86
Qwen2-VL-72B-Instruct [70] Alibaba $\times$ 47.20 0.00 13.67 52.94 45.20 50.33 8.11 53.58
Qwen2.5-VL-72B-Instruct [5] Alibaba $\times$ 55.60 1.20 11.27 49.13 41.20 45.30 11.77 57.04
Qwen-Omni-Turbo [67] Alibaba $\times$ 28.00 0.00 7.97 30.89 14.00 22.84 5.52 43.44
GLM-4V-Plus [64] Zhipu AI $\times$ 54.40 0.00 10.53 51.57 32.80 37.46 8.08 50.38
SkyWork-R1V1-38B [50] Skywork 36.40 0.40 7.91 38.49 3.60 12.55 0.27 29.34
Table 3: The performace evaluations of advanced MLLMs on jigsaw puzzle.
Method Organization Reason- oriented Stage-1 Stage-2 Stage-3
ACC ACC Recall_o Legality ACC Recall_o ThinkGain Legality
Commercial proprietary MLLMs
GPT-4o-20240513 [47] OpenAI $\times$ 26.40 2.00 28.60 100 2.00 31.42 0.85 86.90
o3 [48] OpenAI 34.80 18.00 53.27 98.80 24.00 58.22 1.38 93.93
o4-mini [49] OpenAI 32.40 0.80 16.00 88.40 1.60 12.49 -0.32 35.98
Gemini-2.5-pro-preview-03-25 [63] Google 38.40 20.80 57.00 100 28.40 64.09 0.13 97.17
Gemini-2.0-flash [62] Google 11.60 5.60 41.67 99.93 10.80 39.17 -0.53 72.13
Claude-3.5-Sonnet-20241022 [3] Anthropic 34.40 0.80 26.60 99.87 0.80 9.34 -1.55 23.89
Step-1o-vision-32k [58] StepFun $\times$ 32.20 0.00 20.87 98.87 1.20 22.93 -0.01 92.70
Doubao-1.5-vision-pro-32k [7] ByteDance $\times$ 22.80 0.00 16.73 100 0.00 16.10 0.07 88.97
SenseChat-Vision-V5.5 [54] SenseTime $\times$ 26.40 1.60 32.87 98.80 1.60 31.91 -0.26 97.26
Hunyuan-Vision-20250103 [68] Tencent $\times$ 24.80 0.00 16.80 98.80 0.00 12.10 -1.30 69.63
Open-source MLLMs
LLaVA-OneVision-72B [32] ByteDance & NTU $\times$ 14.40 0.00 17.27 100 0.00 15.80 0.00 100
InternVL-2.5-78B [12] Shanghai AI Lab $\times$ 27.20 2.00 31.73 100 2.40 28.25 -0.27 89.24
InternVL-3-78B [87] Shanghai AI Lab $\times$ 24.80 0.80 26.60 100 4.40 33.24 -0.20 84.85
Qwen2-VL-72B-Instruct [70] Alibaba $\times$ 25.20 0.40 19.47 100 0.40 19.48 0.06 99.61
Qwen2.5-VL-72B-Instruct [5] Alibaba $\times$ 34.00 0.40 25.13 99.93 0.80 25.75 -0.33 94.25
Qwen-Omni-Turbo [67] Alibaba $\times$ 20.80 0.00 17.33 99.60 0.00 16.88 -0.14 90.62
GLM-4V-Plus [64] Zhipu AI $\times$ 34.80 0.40 22.07 100 0.80 14.36 0.00 61.74
SkyWork-R1V1-38B [50] Skywork 20.40 0.00 15.13 80.80 0.00 14.60 0.00 74.89
5.1 Experimental Setup
Evaluation MLLMs. We select 18 top-performing MLLMs for comprehensive evaluation. Regarding proprietary models, we include powerful OpenAI and Gemini series, such as o3/o4 [49] and Gemini-2.5-pro [63]. We also test popular models from the competitive organizations, such as Claude-3.5-Sonnet [3] and Doubao [7]. As for open-source models, we select the respresentive series, such as InternVL and Qwen-VL [87, 5, 64]. Besides, we select the varying versions of a model family with its largest capacity to investigate performance gaps. Finally, we evaluate MLLMs with targeted thinking capabilities, such as SkyWork-R1V1-38B [50]. Note that we take GPT-4o [47] as the baseline model.
Implementation Details. We access proprietary and open models via APIs and local deployment. The maximum token limit is 8192, temperature is 0, both Top-K and Top-P are 1. The rest of hyper-parameter settings keeps same with VLMEvalKit [20]. All the experiments are conducted on a machine with 8 $\times$ NVIDIA A100 GPUs (80G). Due to space constraint, we only apply IPII strategy for maze navigation here. Specifically, Level-1 prompts are consistent with origin prompts, which are illustrated in Fig. 7. Level-2 builds upon Level-1 by incorporating the following prompting information: Please make sure that after executing the move at each step, you should update your internal intermediate visual state, rather than remaining in the initial input visual state, as displayed in Fig. 11. As shown in Fig. 12, Level-3 further builds upon Level-2 by adding the coordinates information of the starting and end points in the maze navigation task.
Table 4: The performace evaluations of advanced MLLMs on embodied long-horizon planning.
Method Organization Reason- oriented Stage-1 Stage-2
ACC ACC Recall_o
Commercial proprietary MLLMs
GPT-4o-20240513 [47] OpenAI $\times$ 55.60 0.80 2.68
o3 [48] OpenAI 62.80 1.60 2.52
o4-mini [49] OpenAI 62.00 0.40 2.25
Gemini-2.5-pro-preview-03-25 [63] Google 60.80 18.80 25.21
Gemini-2.0-flash [62] Google 42.80 4.80 9.06
Claude-3.5-Sonnet-20241022 [3] Anthropic 54.80 0.40 1.48
Step-1o-vision-32k [58] StepFun $\times$ 66.80 4.00 5.64
Doubao-1.5-vision-pro-32k [7] ByteDance $\times$ 41.60 1.60 1.80
SenseChat-Vision-V5.5 [54] SenseTime $\times$ 30.80 0.80 1.72
Hunyuan-Vision-20250103 [68] Tencent $\times$ 30.40 2.80 7.96
Open-source MLLMs
LLaVA-OneVision-72B [32] ByteDance & NTU $\times$ 50.00 2.40 3.54
InternVL-2.5-78B [12] Shanghai AI Lab $\times$ 11.60 0.40 0.47
InternVL-3-78B [87] Shanghai AI Lab $\times$ 29.20 2.00 2.98
Qwen2-VL-72B-Instruct [70] Alibaba $\times$ 53.20 1.60 2.99
Qwen2.5-VL-72B-Instruct [5] Alibaba $\times$ 61.20 1.20 2.61
Qwen-Omni-Turbo [67] Alibaba $\times$ 40.40 0.80 1.29
GLM-4V-Plus [64] Zhipu AI $\times$ 47.20 5.60 9.05
SkyWork-R1V1-38B [50] Skywork 27.20 0.40 1.26
Table 5: The performace evaluations of advanced MLLMs on complex counting.
Method Organization Reason- oriented Stage-1 Stage-2 Stage-3
ACC ACC Recall_o ACC Recall_o ThinkGain
Commercial proprietary MLLMs
GPT-4o-20240513 [47] OpenAI $\times$ 28.80 12.00 44.30 14.40 41.50 7.60
o3 [48] OpenAI 38.40 10.40 40.00 17.60 50.70 10.10
o4-mini [49] OpenAI 40.00 2.40 25.40 3.20 24.70 6.20
Gemini-2.5-pro-preview-03-25 [63] Google 41.20 18.80 36.60 18.00 43.60 0.00
Gemini-2.0-flash [62] Google 46.80 5.60 22.00 12.80 33.50 1.40
Claude-3.5-Sonnet-20241022 [3] Anthropic 21.60 3.20 22.00 4.40 28.90 6.80
Step-1o-vision-32k [58] StepFun $\times$ 43.20 9.60 28.50 6.80 27.80 0.00
Doubao-1.5-vision-pro-32k [7] ByteDance $\times$ 43.60 13.20 33.20 6.00 21.80 -11.00
SenseChat-Vision-V5.5 [54] SenseTime $\times$ 24.00 14.80 44.70 15.20 44.10 5.10
Hunyuan-Vision-20250103 [68] Tencent $\times$ 32.00 8.00 30.90 3.20 16.00 -0.30
Open-source MLLMs
LLaVA-OneVision-72B [32] ByteDance & NTU $\times$ 35.20 6.00 22.60 8.00 24.20 -1.70
InternVL-2.5-78B [12] Shanghai AI Lab $\times$ 21.20 6.40 23.00 4.80 22.70 0.30
InternVL-3-78B [87] Shanghai AI Lab $\times$ 28.80 11.60 34.90 15.20 37.00 0.00
Qwen2-VL-72B-Instruct [70] Alibaba $\times$ 36.00 11.60 36.20 12.40 33.00 -0.60
Qwen2.5-VL-72B-Instruct [5] Alibaba $\times$ 50.00 8.00 32.90 10.00 29.70 0.00
Qwen-Omni-Turbo [67] Alibaba $\times$ 27.60 16.40 44.00 10.00 26.60 -7.90
GLM-4V-Plus [64] Zhipu AI $\times$ 37.20 14.40 43.00 9.20 37.10 0.00
SkyWork-R1V1-38B [50] Skywork 36.80 6.40 27.30 7.60 26.70 0.00
5.2 Main Results
Maze Navigation. Tab. 2 displays that most MLLMs exhibit competent performance in Stage 1. Performance significantly drops in Stage 2, indicating that current MLLMs have limitations in open-ended spatial reasoning and perception. In Stage 3, with the supports of free-style VIS, all models consistently achieves gains in global-level ACC and fine-grained $\mathbf{R\_o}$, leading to impressive ThinkGain, which indicates the effectiveness of free-style IVS in tackling deficiencies of spatial-aware cognition. However, we observe decline in Legality, which indicates that external knowledge could further confuse the thinking trajectories of weaker models. From Fig. 4, we find that response length is inversely proportional to efficiency in maze navigation. These outcomes suggest that free-style IVS can enhance performance in spatial reasoning for MLLMs with strong learned reasoning priors.
Jigsaw Puzzle. Tab. 3 indicates that there also exists significant declines from Stage 1 to 2, confirming that open-ended QA poses challenges to visual CoT in this task, which is likely to stem from the difficulties in understanding AIGC semantics and distinguishing semantically incoherent patches. MLLMs in Stage 3 promoted by free-style IVS exhibit global-level gains compared to Stage 2, but surprisingly achive negative ThinkGain with impressive drops in Legality, which could be due to the irrational IVS utiliztions leading to invalid decisions. From Fig. 4, we figure out that top-performing models [48, 63] have concise responses, while SkyWork-R1V1 [50] with longer responses exhibits wipeout. These failures shows the deficiencies of current models in using free-form IVS, reflecting the inadequate VI-CoT. Future works should explore promoting positive utilizations of IVS in training.
Figure 4: Quantitative comparisons on averge response length, ACC, and ThinkGain for three tasks.
Figure 5: The further investigations with IPII strategy in VI-CoT capability for maze navigation.
Embodied Long-Horizon Planning. Most models exhibit significant declines from Stage 1 to 2, particularly Step-1o [58]. We attribute this observation to models exploiting multiple-choice QA by matching semantics with the given sequence to make hacking decision. However, results in Stage 2 demonstrate that models do not truly understand the instructions and engage in reasoning. Instead, they tend to directly make plans based on semblances. For Stage 3, we select the real-world scenario for qualitative comparison, as shown in Fig. 6. We conclude that powerful models can grasp implicit instructions to perform reasoning for object searching, indicating the benefits of free-style IVS in facilitating embodied planning.
Complex Counting. ACC in Stage 2 is inferior to Stage 1, which confirms our conclusions in above tasks. In Stage 3, some models [7, 67, 68] exhibit ACC declines and negative ThinkGain. Surprisingly, Gemini-2.5-pro [63] also exhibits drops with null ThinkGain. These outcomes indicate that MLLMs still suffer from object hallucinations and deficiencies in basic perception, leading to accumulated erroneous trajectories. In case studies, we also find that IVS assist models in recognizing heads, yet they still make wrong decisions, consistent with observations in LLMs [22, 78]. From Fig. 4, we find that models with relatively better performance tend to have shorter responses. The above results reveal the impressive difficulties MLLMs face with counting, which should be promoted by developing targeted representations or training strategies.
Further Investigations with IPII Strategy. Proprietary models in Stage 1 consistently achieve better performance. Among open models, [87, 5] experience varying degrees of decline. Surprisingly, SkyWork-R1V1 [50] finally reaches 98.40% ACC, indicating that R1-like MLLMs might have higher sensitivities to prompts. In Stage 2, all models do not achieve impressive outcomes across levels, revealing that existing MLLMs might lack the capability to implicitly update their internal IVS through textual CoT in the open-ended QA. These results also indicates the necessity of utilizing free-style IVS as the external knowledge to improve reasoning. In Stage 3, performance is closely related to the inherent VI-CoT priors. Strong models [63] can leverage external knowledge to achieve better results, while weaker models might be confused by extra knowledge, leading to drops.
Figure 6: The qualitative analysis for Stage 3 with free-style IVS in embodied long-horizon planning.
6 Conclusion and Limitations
We introduce ViC-Bench, a specialized benchmark designed to evaluate VI-CoT capability in MLLMs. Our dataset consists of four representative tasks, with each has dedicated construction and free-style IVS generation pipelines supporting function calls. To obtain a thorough understanding of VI-CoT performance, we design a novel progressive three-stage evaluation suite with targeted new metrics. The IPII strategy also impressively indicate the prompting factors which affect VI-CoT performance. The systematic evaluations obtain key observations and insights into the current developments of VI-CoT in MLLMs. We hope ViC-Bench can inspire more research in multi-modal reasoning. As the field continues to evolve, ViC-Bench stands as a valuable tool for measuring progress in the development of more sophisticated multi-modal systems.
Despite our efforts, limitations still exist. The ThinkGain metric involves black-boxing the retrospection of MLLMs, and we plan to deeply delve into retrospection for more detailed investigations. Besides, we will supplement the quantitative results of Stage 3 for embodied long-horizon planning in the future.
7 Acknowledgements
Our ViC-Bench follows the licenses of the public datasets involved in our construction phase and will not be transferred to outside parties without permissions and can be only utilized for academic research. In particular, this benchmark dataset will not be included as a part of any commercial software package or product of any institution.
Appendix A Appendix
Global-Aware Prompts for Four Downstream Tasks
To more intuitively conduct the comprehensive comparisons of MLLMs performance in VI-CoT capability, Figs. 7, 8, 9, & 10 below intimately display the global-aware prompts for maze navigation, jigsaw puzzle, embodied long-horizon planning, as well as complex counting tasks across three-stage evaluations. Besides, we also present the prompting-style instructions in Stage 3 of each downstream task that couple with the free-style IVS when delivered under the function call paradigm.
Weighted Dispersed Sampling Strategy
Algorithm 1 Weighted Dispersed Sampling Strategy
1:Grid size $G$, Number of rows $R$
2:Sorted list $S_{final}$ of $N=6$ selected patch indices
3:$S\leftarrow\emptyset$triangleright Initialize set of selected indices
4:$N\leftarrow 6$triangleright Target number of patches
5:$C\leftarrow G/R$triangleright Patches per row (assuming integer division)
6:if $C\leq 0$ or $R\leq 0$ or $G<R$ then
7:     raise ValueError("Invalid grid/row configuration")
8:Let $I_{r}=\{i\mid r\cdot C\leq i<(r+1)\cdot C\}$ for $r=0,\dots,R-1$triangleright Indices for each row
9:$S_{init}\leftarrow\emptyset$
10:for $r\leftarrow 0$ to $R-1$ do
11:     if $I_{r}\neq\emptyset$ then
12:         Choose $i_{r}$ uniformly at random from $I_{r}$
13:         $S_{init}\leftarrow S_{init}\cup\{i_{r}\}$
14:$S\leftarrow\text{Unique}(S_{init})$triangleright Ensure uniqueness
15:while $|S|<R$ and $|S|<G$ do $\triangleright$ Fill up to R indices if needed
16:     $R_{avail}\leftarrow\{0,\dots,G-1\}\setminus S$
17:     if $R_{avail}=\emptyset$ then
18:         break
19:     Choose $i_{new}$ uniformly at random from $R_{avail}$
20:     $S\leftarrow S\cup\{i_{new}\}$
21:while $|S|<N$ do
22:     $R_{avail}\leftarrow\{0,\dots,G-1\}\setminus S$triangleright Remaining available indices
23:     if $R_{avail}=\emptyset$ then
24:         break
25:     $W_{all}\leftarrow\text{calculate\_weights}(G,S)$triangleright Get weights for all grid cells based on current $S$
26:     $W_{rem}\leftarrow\{W_{all}[j]\mid j\in R_{avail}\}$triangleright Extract weights for available indices
27:     $W_{sum}\leftarrow\sum_{w\in W_{rem}}\max(w,10^{-9})$triangleright Sum of remaining weights (avoiding zero)
28:     if $W_{sum}>10^{-9}$ and $|R_{avail}|>0$ then
29:         $P\leftarrow[\max(W_{all}[j],10^{-9})/W_{sum}\text{ for }j\in R_{avail}]$triangleright Calculate probabilities
30:         $P\leftarrow P/\sum P$triangleright Normalize probabilities rigorously
31:     else if $|R_{avail}|>0$ then
32:         $P\leftarrow[1/|R_{avail}|\text{ for }j\in R_{avail}]$triangleright Fallback to uniform probability
33:     else
34:         break
35:     try
36:     Choose $i_{new}$ from $R_{avail}$ according to probability distribution $P$
37:     $S\leftarrow S\cup\{i_{new}\}$
38:     catch ValueError $\triangleright$ If probability choice fails
39:     if $R_{avail}\neq\emptyset$ then
40:         Choose $i_{new}$ uniformly at random from $R_{avail}$
41:         $S\leftarrow S\cup\{i_{new}\}$
42:     else
43:         break
44:     end try
45:$S_{final}\leftarrow S$
46:if $|S_{final}|>N$ then
47:     $S_{final}\leftarrow\text{RandomSample}(S_{final},N)$triangleright Sample down to N
48:else if $|S_{final}|<N$ then
49:     $R_{final}\leftarrow\{0,\dots,G-1\}\setminus S_{final}$
50:     $N_{needed}\leftarrow N-|S_{final}|$
51:     $N_{add}\leftarrow\min(N_{needed},|R_{final}|)$
52:     if $N_{add}>0$ then
53:         $S_{extra}\leftarrow\text{RandomSample}(R_{final},N_{add})$
54:         $S_{final}\leftarrow S_{final}\cup S_{extra}$
55:if $|S_{final}|\neq N$ then
56:     Print warning "Final selected count is $|S_{final}|$, target was $N$"
57:return $\text{SortAscending}(S_{final})$
Helper Functions we deployed:
  • $\bullet$
[…]
  • $\bullet$
    $\text{Unique}(S)$: Returns a set containing unique elements from list/set $S$.
  • $\bullet$
[…]
  • $\bullet$
[…]
  • $\bullet$
    $\text{RandomSample}(L,k)$: Selects $k$ unique elements uniformly at random from list $L$.
  • $\bullet$
    $\text{SortAscending}(L)$: Returns a new list containing elements of $L$ sorted in ascending order.
Figure 7: The overall illustrations of three-stage prompts for maze navigation task.
Figure 8: The overall illustrations of three-stage prompts for jigsaw puzzle task.
Figure 9: The overall illustrations of three-stage prompts for embodied long-horizon planning task.
Figure 10: The overall illustrations of three-stage prompts for complex counting task.
Figure 11: The overall illustrations of the Level-2 prompts in IPII strategy for maze navigation task.
Figure 12: The overall illustrations of the Level-3 prompts in IPII strategy for maze navigation task.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We have clearly stated our contributions in both Abstract and Introduction sections, including the proposal of a new benchmark denoted ViC-Bench and the new evaluation suite.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of our work are discussed as thoroughly as possible in Sec. 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We conduct detailed evaluations with our introduced ViC-Bench and presents extensive experimental results. We also perform ablation prompting studies to validate the effectiveness of our theoretical findings.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We have provided a clear description of the specific details of our benchmark construction and three-stage evaluation suite. Besides, we will make our benchmark and evaluation code publicly available for the reproducibility.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: During review, we have provided our data and related code with reviewers and chairs. Moreover, we will make data and code publicly open after the paper decision.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We specify all the implementation details in Sec. 5.1.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: We do not conduct any experiment on the statistical significance.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The details of our experiments compute resources are described in Sec. 5.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: : Our constructed benchmark incorporates the ethical considerations, as explicitly stated in the paper.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: In Sec. Introduction and Conclusion, we have discussed the potential scenarios of our established benchmark in the research community of MLLMs.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: Considering the potential for malicious use of our proposed ViC-Bench, we have included user guidelines for ViC-Bench in the Appendix.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have endeavored to include URLs wherever possible for referenced content, and proper acknowledgment has been given to the creators of the assets used in this paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: Throughout this paper, we have offered the detailed descriptions of the use of MLLMs or LLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.